OpenStudy (anonymous):

if the points (3,r) and (r,-1) lie on a line with slope (-3/4) what is the value of r?

OpenStudy (owlcoffee):

okay so we have got 2 points... Let's call them A and B. where A(3,r) and B(r,-1).

OpenStudy (owlcoffee):

we want to find the value of r, so let's tke those points and find the equation of the line that goes through those points.

OpenStudy (owlcoffee):

We have the slope, so it would be optimal to do this exercise with the point-slope formula and taking arbitrary A or B. Do you follow?

OpenStudy (anonymous):

I understand all except the last response, can you explain what arbitrary means?

OpenStudy (owlcoffee):

arbitrary means basically "any you want" so you can choose any of those points to apply the point-slope formula.

OpenStudy (anonymous):

so where would we go from there?

OpenStudy (owlcoffee):

Now, we have two points, i'll pick point A(3,r) and use the point slope formula: \[y-y _{1}=m(x-x_{1})\] where m is the slope and x1, y1 the point's coordinates. I'll use that for A(3,r), ending up with: \[y-r=-\frac{ 3 }{ 4 }(x-3)\] I'll apply distributive to get rid if the -3/4: \[y-r=-\frac{ 3x }{ 4 }+\frac{ 9 }{ 4 }\] continuing let's add "r" to both sides and use common denominator: \[y=\frac{ -3x+9+4r }{ 4 }\] multiplying both sides by 4 we get: \[4y=-3x+9+4r\] So now we found an equation that can help us find r, and we know that point B belongs to this line, so we can replace x and y with the coordinates of B: \[-4=-3r+9+4r\] operating a little, we get: \[(4-3)r=9-4\] and our final answer is: \[r=5\] Not hard, was it. Now try doing it but picking B first.

OpenStudy (anonymous):

That makes sense, thank you! but I don't think i did point b right r+1=(-3/4)(3-r) r+1=-9/4-3/4r 1/4r=-13/4 r=-13

OpenStudy (owlcoffee):

oh, you have to write the point-slope equation for just one point, simplify and then replace the values of the points.

OpenStudy (anonymous):

so is the value of r 5 or -13?

OpenStudy (anonymous):

5 right? i think i see where i went wrong

OpenStudy (anonymous):

yeah i set the equation up wrong, thank you so much!!!

OpenStudy (owlcoffee):

No problem, I love helping people.